OpenStudy (anonymous):

solving for x is 28 shapes are similar/same just different measures

OpenStudy (anonymous):



OpenStudy (anonymous):

8/6 = x/21 x = 21(8)/6 x = 168 / 6 x = 28

OpenStudy (anonymous):

thanks